OpenStudy (anonymous):

Vectors Prove that (a x b) . (c + xa + yb) = a.(b x c) for any numbers x and y.

OpenStudy (turingtest):

prove\[\forall x,y\in\mathbb R:(\vec a\times\vec b)\cdot(\vec c+x\vec a+y\vec b)=\vec a\cdot(\vec b\times \vec c)\]is this the problem?

OpenStudy (anonymous):

Yes. Sorry, I struggled with the notation.

OpenStudy (anonymous):

im thinking that \(\vec a \times \vec b\) is perpendicular to \(\vec a\) and \(\vec b\)

OpenStudy (turingtest):

good call :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

....forgive me, you may need to explain this to me...

OpenStudy (anonymous):

note that after expanding we have\[(\vec a\times\vec b)\cdot(\vec c+x\vec a+y\vec b)=(\vec a\times\vec b). \vec c+x(\vec a\times\vec b).\vec a+y(\vec a\times\vec b).\vec b\]

OpenStudy (turingtest):

...and what is the product of\[(\vec a\times\vec b)\cdot\vec a\]?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

Ah. So you're left with the triple scalar product abc. Thanks